If two resistors R1 and R2 are connected in parallel, then the total resistance R measured in ohms is given by

1/R=1/R1+1/R2
If the resistance of R1 and R2 are increasing at 390 and 670 ohms per second respectively, then the rate of increase of R when R1=7000 and
What is R2=3000 ohms per second?

Can you show me the work? I will calculate the answer. Thanx!
Answer #1
1/R=1/R1+1/R2
differentiating
1/R^2 dR/dt=1/R1^2 dR1/dt + 1/R2^2 dR2/dt
R=390*670/(670+390)=246.5
dR1/dt=7000
dR2/dt=3000
R1=390
R2=670
putting in equation
we get dR/dt=3202.496 ohms per second
answered by: donald tegue